# Federated Unlearning: a Perspective of Stability and Fairness

### v1

@@ Abstract
This paper explores the multifaceted consequences of federated unlearning (FU) with data heterogeneity. We introduce key metrics for FU assessment, concentrating on verification, global stability, and local fairness, and investigate the inherent trade-offs. Furthermore, we formulate the unlearning process with data heterogeneity through an optimization framework. Our key contribution lies in a comprehensive theoretical analysis of the trade-offs in FU and provides insights into data heterogeneity’s impacts on FU. Leveraging these insights, we propose FU mechanisms to manage the trade-offs, guiding further development for FU mechanisms. We empirically validate that our FU mechanisms effectively balance trade-offs, confirming insights derived from our theoretical analysis.
@@ 1 Introduction
With the advancement of user data regulations, such as GDPR Regulation (2018) and CCPA Goldman (2020), the concept of “the right to be forgotten” has gained prominence. It necessitates models’ capability to forget or remove specific training data upon users’ request, which is non-trivial since the models potentially memorize training data. Intuitively, the most straightforward approach is to exactly retrain the model from scratch without the data to be forgotten. However, this method is computationally expensive, especially for large-scale models prevalent in modern applications. As a result, the machine unlearning (MU) paradigm is proposed to efficiently remove data influences from models Cao & Yang (2015). The effectiveness of unlearning, measured by verification approaches, requires the unlearning mechanism to closely replicate the results of exact retraining without heavy computational burden.
Federated learning (FL) has gained attention in academia and industry with increased data privacy concerns by allowing distributed clients to collaboratively train a model while keeping the data local Kairouz et al. (2021). While MU offers strategies for traditional centralized machine learning context, federated unlearning (FU) introduces new challenges due to inherent data heterogeneity and privacy concerns in the federated context Wang et al. (2023). Recent research of FU, such as Gao et al. (2022); Che et al. (2023), mainly focused on verification and efficiency in FU, aligning with the main objectives of MU. However, the inherent data heterogeneity in federated systems introduces new challenges: (i) due to clients’ diverse preferences for the global model, unlearning certain clients could result in unequal impacts on individuals; (ii) different clients contribute differently to the global model, thus unlearning specific clients can lead to diverse impacts on model performance.
Figure 1: Federated Unlearning and Its Consequences. (a) We consider an FL system of $5$ clients with non-IID training data. One client requests to be unlearned. (b) We conduct $5$ experiments, each featuring the unlearning of one specific client.
On the challenges of FU under heterogeneous data.
Figure 1 elaborates two key insights into the challenges posed by data heterogeneity in FU. (1) Local Fairness: As shown in Figure 1(a), FU can unequally impact remaining clients, where some clients benefit from unlearning, but others experience disadvantages. It illustrates a “local fairness” concern, pertaining to the uniform distribution of utility changes [^1] In this context, ‘utility’ refers to an individual client’s experienced performance of the global model. among remaining clients after unlearning. (2) Global Stability: Unlearning different clients leads to different impacts on the global model’s performance, as depicted in Figure 1(b). This highlights the “global stability” concern in FU, emphasizing the need to maintain consistent system performance. These empirical findings identify two trade-offs inherent in FU: the FU verification vs. global stability trade-off, and the FU verification vs. local fairness trade-off. These insights motivate us to ask:
How can we assess the consequences of FU, and what are the theoretical and practical approaches to balancing the inherent trade-offs?
@@
To address the above question, we will construct a comprehensive theoretical framework for FU’s consequences, which should offer a rigorous understanding of how data heterogeneity affects FU while balancing the trade-offs.
Contributions. We outline the structure of our paper in Figure 2 and summarize our key contributions as follows:
@@
    Qualitative Understanding of FU Metrics: We introduce robust quantitative metrics for FU assessment, including FU verification metric, global stability metric and local fairness metric (detailed in Section 4). These metrics provide a foundation for comprehensive trade-off evaluations in FU.
@@
    FU Mechanism and Theoretical Framework: We propose a novel FU mechanism based on the theoretical framework, encompassing optimization strategies and penalty methods. We provide the theoretical analysis and practical insights into balancing the tradeoffs within a verifiable FU context, as detailed in Section 6.
  4. 4.
    Empirical Validation: In Section 7, we empirically validate our FU mechanisms in non-convex settings, confirming theoretical insights by effectively balancing trade-offs.
@@ 2 Related Work
Mechine Unlearning & Federated Unlearning. Machine unlearning (MU) aims to remove specific data from a machine learning model, addressing the challenges in both effectiveness and efficiency of the unlearning process Guo et al. (2020); Wu et al. (2020); Bourtoule et al. (2021b; a); Tarun et al. (2023a; b); Jia et al. (2023). In FL, federated unlearning (FU) is proposed to address clients’ right to be forgotten, including methods like rapid retraining Liu et al. (2022), subtracting historical updates from the trained model Wu et al. (2022), subtracting calibrated gradients of the unlearn clients to remove their influence Liu et al. (2020; 2021), and adding calibrated noises to the trained model by differential privcacy Zhang et al. (2023). However, none of them involves any rigorous consideration for data heterogeneity, the main challenge in FL. In this work, we account for data heterogeneity in FU through a comprehensive optimization framework, and theoretically analyze how data heterogeneity impacts unlearning in Section 5.
@@
Moreover, existing FU methods has focused on methods ensuring verifiable and efficient unlearning Liu et al. (2021; 2022); Fraboni et al. (2022); Gao et al. (2022); Jin et al. (2023); Che et al. (2023). The verification of unlearning typically involves comparing the unlearned model, obtained through an unlearning mechanism, with a reference model using performance metrics such as accuracy and model similarity metrics Gao et al. (2022). Additionally, attack-based verification methods, such as membership inference attacks (MIA) and backdoor attacks (BA), are often employed in MU Nguyen et al. (2022) but are not applicable in federated systems due to privacy concerns. Besides verification, data heterogeneity in FU introduces consequences on global stability and local fairness, necessitating consideration of trade-offs in FU, as previously discussed in Section 1. In this work, we conduct a rigorous analysis of inherent trade-offs in FU under data heterogeneity .
@@
Stability in FL and FU. In FL, performance stability revolves around maintaining consistent and robust model performance despite alterations in the training dataset. This aspect of stability, highlighted in studies such as Yin et al. (2018); Fang et al. (2020); Li et al. (2021), often involves defending against external threats. Unlike FL, FU is concerned with managing internal changes within the system for users’ rights to be forgotten. Considering unlearning in the federated system, the inherent data heterogeneity can lead to significant shifts in data distribution, thereby altering the model’s performance. In this work, we delve into the theoretical analysis of unlearning, examining how data heterogeneity impacts stability in FU.
@@
Fairness in FL and FU. In FL, there are several works that have proposed different notions of fairness. The proportional fairness ensures whoever contributes more to the model can gain greater benefits Wang et al. (2020); Yu et al. (2020). Additionally, the model fairness focuses on protecting specific characteristics, like race and gender Gu et al. (2022). Furthermore, the performance fairness Li et al. (2019); Mohri et al. (2019); Hao et al. (2021); Li et al. (2021); Shi et al. (2023); Pan et al. (2023) aims to reduce the variance of local test performance or utility across all clients.
In FU, we observe that unlearning certain clients can lead to unequal impacts on remaining clients due to data heterogeneity, as discussed in Section 1. In this work, we extend performance fairness to FU by the variance of utility changes among remaining clients and further analyze how data heterogeneity impacts fairness in FU.
@@ 3 Preliminaries
3.1 Federated Learning (FL)
Suppose there is a client set $\mathcal{N}$ ($|\mathcal{N}|=N$), contributing to FL training. Each client $i\in\mathcal{N}$ has a local training dataset with size $n_{i}$, and the data is non-IID across different clients. The optimal FL model is defined below:
Definition 3.1 (Optimal FL Model).
The optimal solution of the FL global model, denoted as $\boldsymbol{w}^{*}$, can be expressed by the following optimization problem:
$\boldsymbol{w}^{*}=\arg\min_{\boldsymbol{w}}[F(\boldsymbol{w}):=\sum_{i\in\mathcal{N}}p_{i}f_{i}(\boldsymbol{w})]$
Here, $f_{i}$ denotes the local objective function of client $i$ with the aggregation weight $p_{i}=\frac{n_{i}}{\sum_{k\in\mathcal{N}}n_{k}}$.
The empirically trained model $\boldsymbol{\boldsymbol{w}}^{o}$ is obtained by training all clients $i\in\mathcal{N}$, serving as the empirical approximation of the theoretical optimum $\boldsymbol{w}^{*}$ (summarized in Figure 3).
3.2 Federated Unlearning (FU)
FU mechanisms aim to obtain an unlearned model $\boldsymbol{{w}}^{u}$ that removes the influence of client set $\mathcal{J}$ who requests to be forgotten from the original trained model $\boldsymbol{\boldsymbol{w}}^{o}$. This paper focuses on the FU mechanism that begins with $\boldsymbol{\boldsymbol{w}}^{o}$ and iteratively updates it through the participation of the remaining clients ($i\notin\mathcal{J}$). The model is refined over $T$ rounds akin to FL, converging to the unlearned model $\boldsymbol{{w}}^{u}$. The optimal unlearned model could be defined as:
Definition 3.2 (Optimal FU Model).
In the context of FU, the optimal unlearned model is defined as the minimizer for the global objective $F_{-\mathcal{J}}$ of all remaining clients $i\notin\mathcal{J}$. Formally, it is expressed as:
$\boldsymbol{w}^{r*}=\arg\min_{\boldsymbol{w}}[F_{-\mathcal{J}}(\boldsymbol{w}):=\sum_{i\notin\mathcal{J}}p_{i}^{\prime}f_{i}(\boldsymbol{w})],$
Here, $p_{i}^{\prime}$ is the normalized aggregation weight during unlearning, i.e., $p_{i}^{\prime}=\frac{p_{i}}{1-P_{\mathcal{J}}}$ with $P_{\mathcal{J}}=\sum_{j\in\mathcal{J}}p_{j}$.
The exact retrained model $\boldsymbol{w}^{r}$ is obtained after retraining all remaining clients [^2] In this paper, the retraining employs FedAvg McMahan et al. (2017)., serving as the empirical approximation of the theoretical optimum $\boldsymbol{w}^{r*}$ (summarized in Figure 3).
Figure 3: Key Model Notations ($\boldsymbol{{w}}^{u}$ represents the unlearned model from the FU mechanism; $\boldsymbol{w}^{r}$ is from retraining with remaining clients, the special case of $\boldsymbol{{w}}^{u}$).
@@ 4.1 FU Verification
Verification is critical to evaluate how much the unlearning mechanism effectively removes the data Yang & Zhao (2023). Previous studies on unlearning verification often utilize a weak reference like $\boldsymbol{w}^{r}$ from retraining for comparison Halimi et al. (2022); Che et al. (2023), but this reference is often inconsistent for variability and randomness in the retraining process. Thus, we employ the optimal unlearned model $\boldsymbol{w}^{r*}$ as a theoretical benchmark for verification, allowing for consistent and replicable evaluations in FU. We define a verification metric by the performance gap between the unlearned model and $\boldsymbol{w}^{r*}$:
@@
where $F_{-\mathcal{J}}(\cdot)$ measures performance over the remaining clients after unlearning certain clients set $\mathcal{J}$.
@@ 4.2 Global Stability
Effective unlearning often requires modification to the trained model, which can lead to performance variations of the global model. In FL, maintaining stable performance is particularly challenging under data heterogeneity, as unlearning critical clients can significantly compromise model performance. To measure the extent of performance stability in FU, we propose the global stability metric as follows:
@@ 4.3 Local Fairness
In FL, fairness can be associated with the consistency of model performance across different clients. Specifically, a model is considered fairer if its performance has a smaller variance across clients Li et al. (2020a).
@@
where $\Delta f_{i}(\boldsymbol{{w}}^{u})=\mathbb{E}\left[f_{i}(\boldsymbol{{w}}^{u})\right]-f_{i}(\boldsymbol{w}^{*})$ represents the utility change for remaining client $i\notin\mathcal{J}$ due to FU. Moreover, $\overline{\Delta f}=\sum_{i\notin\mathcal{J}}p_{i}^{\prime}\Delta f_{i}(\boldsymbol{{w}}^{u})$ is the weighted average of local utility changes among remaining clients, serving as a benchmark for assessing deviation in the impacts of unlearning.
The metric $Q$ is inspired by the mean absolute deviation (MAD) in measuring fairness of FL Ezzeldin et al. (2023). The metric $Q$ captures FU’s impact on utility changes experienced by remaining clients and further facilitates theoretical analysis of fairness implication in Section 5.2.
@@ 5 Theoretical Analysis on Trade-offs in FU
This section provides a theoretical analysis of the trade-offs in FU, particularly focusing on the balance between FU verification and stability, as well as FU verification and fairness, as outlined in Section 5.1 and 5.2, respectively. Our analysis critically examines the challenges posed by data heterogeneity in FU. To begin, we formally state the assumptions required for the theoretical analysis.
@@
Let the objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be $\mu$-strong convex. For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, the function $f$ satisfies the following inequality: $f(\mathbf{u})\geq f(\mathbf{v})+\langle\nabla f(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle+\frac{\mu}{2}\|\mathbf{u}-\mathbf{v}\|^{2}$, where $\mu>0$ is the convexity constant.
@@
Assume that the objective function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L$-smooth. For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, the function $f$ satisfies the following inequality: $f(\mathbf{u})\leq f(\mathbf{v})+\langle\nabla f(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle+\frac{L}{2}\|\mathbf{u}-\mathbf{v}\|^{2}$, where $L>0$ is the Lipschitz constant of the gradient of $f$.
@@
Given a subset of remaining clients $\mathcal{S}$, the aggregation of their stochastic gradients $\boldsymbol{g}_{\mathcal{S}}(\boldsymbol{w})=\sum_{i\in\mathcal{S}}\alpha_{i}\boldsymbol{g}_{i}(\boldsymbol{w})$ is unbiased estimator of $\nabla F_{-\mathcal{J}}(\boldsymbol{w})$ with bounded variance: $\mathbb{E}\left[\left\|\boldsymbol{g}_{\mathcal{S}}(\boldsymbol{w})-\nabla F_{-\mathcal{J}}(\boldsymbol{w})\right\|^{2}\right]\leq\sigma^{2}$.
@@
5.1-5.4 are commonly used for the FL convergence analysis Li et al. (2020b); Wang et al. (2021). 5.5 assumes unlearned clients’ aggregate weights do not exceed those of the remaining clients. This is crucial to prevent catastrophic consequences, which could undermine the objectives of fairness and stability in FU.

### v4

@@ Abstract
This paper explores the side effects of federated unlearning (FU) with data heterogeneity, concentrating on global stability and local fairness. We introduce key metrics for FU assessment and investigate FU’s inherent trade-offs. Furthermore, we formulate the FU problem through an optimization framework for managing side effects. Our key contribution lies in a comprehensive theoretical analysis of the trade-offs in FU and provides insights into data heterogeneity’s impacts on FU. Leveraging these insights, we propose FU mechanisms to manage the trade-offs, guiding further development for FU mechanisms. We also empirically validate that our FU mechanisms effectively balance trade-offs, confirming insights derived from our theoretical analysis.
@@ 1 Introduction
With the advancement of user data regulations, such as GDPR [41] and CCPA [11], the concept of “the right to be forgotten" has gained prominence. It necessitates models’ capability to forget or remove specific training data upon users’ request, which is non-trivial since the models potentially memorize training data. Intuitively, the most straightforward approach is to exactly retrain the model from scratch without the data to be forgotten. However, this method is computationally expensive, especially for large-scale models prevalent in modern applications. As a result, the machine unlearning (MU) paradigm is proposed to efficiently remove data influences from models [5]. The effectiveness of unlearning, measured by verification approaches, requires the unlearning mechanism to closely replicate the results of exact retraining without heavy computational burden.
Federated learning (FL) has gained attention in academia and industry with increased data privacy concerns by allowing distributed clients to collaboratively train a model while keeping the data local [20]. While MU offers strategies for traditional centralized machine learning context, federated unlearning (FU) introduces new challenges due to inherent data heterogeneity and privacy concerns in the federated context [46]. Recent research of FU, such as Gao et al. [10], Che et al. [6], Pan et al. [39], Liu et al. [34], mainly focused on verification and efficiency in FU, aligning with the main objectives of MU. However, the inherent data heterogeneity in federated systems introduces new challenges: (i) different clients contribute differently to the global model, thus unlearning specific clients can lead to diverse impacts on model performance; (ii) due to clients’ diverse preferences for the global model, unlearning certain clients could result in unequal impacts on individuals.
Figure 1: Federated Unlearning and Its Side Effects. (a) FL system of $5$ clients with non-IID training data. with one client requesting to be unlearned. (b) Unlearning each client through $5$ experiments.
@@
On the Side Effects of FU under heterogeneous data. Figure 1 elaborates two key side effects posed by data heterogeneity. (1) Global Stability: Unlearning different clients leads to different impacts on the global model’s performance, as depicted in Figure 1(b). This highlights the “global stability" concern in system performance; (2) Local Fairness: As shown in Figure 1(a), FU can unequally impact remaining clients, where some benefit from unlearning, but others experience disadvantages. It illustrates a “local fairness" concern, pertaining to the uniform distribution of utility changes [^1] In this context, ‘utility’ refers to an individual client’s experienced performance of the global model. among remaining clients after unlearning.
FU undertakes the potential for problematic unbounded instability and unfairness. Without controlling the instability or unfairness, unlearning may potentially drive all remaining clients to leave the system, leading to catastrophic forgetting [30] and problematic behavior among selfish clients that exploit the resources of others, akin to free-rider attacks [9]. Thus, it is essential to consider the two trade-off inherent in FU, including the FU verification vs. global stability trade-off, as well as the FU verification vs. local fairness trade-off. These insights motivate us to ask:
Q: How can we assess the side effects of FU, and what are the theoretical and practical approaches to balancing the inherent trade-offs?
To address the above question, we construct a comprehensive theoretical framework for FU’s side effects, which offers a rigorous understanding of how data heterogeneity affects FU while balancing the trade-offs. We outline the structure of our paper in Figure 2
Contributions. Our key contributions are summarized as follows:
@@
    Qualitative Understanding of FU Metrics: We introduce robust quantitative metrics for FU assessment, including verification metric, global stability metric and local fairness metric (detailed in Section 4). These metrics provide a foundation for comprehensive FU trade-off evaluations.
@@
    FU Mechanism and Theoretical Framework: We propose a novel FU mechanism based on the theoretical framework, encompassing optimization strategies and penalty methods into balancing the tradeoffs within a verifiable FU context, as detailed in Section 6. The experimental evaluation is provided in Section 7.
@@ 2 Related Work
Mechine Unlearning & Federated Unlearning. Machine unlearning (MU) aims to remove specific data from a machine learning model, addressing the challenges in both effectiveness and efficiency of the unlearning process [13, 51, 3, 4, 44, 45, 18, 22]. In FL, federated unlearning (FU) is proposed to address clients’ right to be forgotten, including methods like rapid retraining [33], subtracting historical updates from the trained model [49], subtracting calibrated gradients of the unlearn clients to remove their influence [31, 32], and adding calibrated noises to the trained model by differential privcacy [56].
@@
Existing FU methods has focused on ensuring verifiable and efficient unlearning [32, 33, 8, 10, 19, 6]. The verification of unlearning typically involves comparing the unlearned model, obtained through an unlearning mechanism, with a reference model using performance metrics such as accuracy and model similarity metrics [10]. Additionally, attack-based verification methods, such as membership inference attacks (MIA) and backdoor attacks (BA), are often employed in MU [37] but are not applicable in federated systems due to privacy concerns. However, none of them involves any rigorous consideration for data heterogeneity, the main challenge in FL. As discussed in Section 1, data heterogeneity can impact the global stability and local fairness of the FL system. In this work, we theoretically analyze how data heterogeneity impacts FU in Section 5.
@@
Stability in FL and FU. In FL, the concept performance stability is introduced for describing robust model performance. This aspect of stability often involves defending against external threats, like alterations in the training dataset [54, 7, 26]. Unlike FL, FU is concerned with managing internal changes within the system for users’ rights to be forgotten. Due to the inherent data heterogeneity of FL system, unlearning can lead to data distribution shifts and model performance change. Besides of reducing the performance change, it is also crucial to ensure unlearning is verifiable in FU. In this work, we theoretically examine how data heterogeneity impacts stability in FU, and how to balance the trade-offs between FU verification and stability.
@@
Fairness in FL and FU. In FL, there are several works that have proposed different notions of fairness. The proportional fairness ensures whoever contributes more to the model can gain greater benefits [48, 55]. Additionally, the model fairness focuses on protecting specific characteristics, like race and gender [12]. Furthermore, the performance fairness [27, 36, 15, 26, 42, 40] aims to reduce the variance of local test performance or utility across all clients. The FL fairness is mainly influenced by data heterogeneity across participating clients. In FU, unlearning certain clients can lead to unequal impacts on remaining clients due to data heterogeneity, as discussed in Section 1. In addition to heterogeneity among remaining clients, we also account for the data heterogeneity between the removed clients and the remaining clients, making fairness analysis more challenging in FU compared to FL. In this work, we establish performance fairness by the variance of utility changes among remaining clients and further analyze how data heterogeneity impacts fairness in FU.
@@ 3 Preliminaries
Figure 3: Key Model Notations ($\boldsymbol{{w}}^{u}$ is the unlearned model from the FU mechanism; $\boldsymbol{w}^{r}$ is exact retrained model, the special case of $\boldsymbol{{w}}^{u}$).
Algorithm 1 Federated Unlearning Mechanism $\mathcal{M}$
1:  Input: $\boldsymbol{\boldsymbol{w}}^{o}$ (original FL trained model), $\mathcal{J}$ (clients to unlearn), $T$ (unlearning rounds)
2:  Initialize $\boldsymbol{w}_{0}\leftarrow\boldsymbol{\boldsymbol{w}}^{o}$
3:  for $t=0$ to $T-1$ do
4:     $\mathcal{S}\leftarrow$ subset of remaining clients $\mathcal{N}\setminus\mathcal{J}$
5:     for client $i\in\mathcal{S}$ in parallel do
6:        $\boldsymbol{w}_{i}^{t+1}\leftarrow\text{ClientUpdate}(\boldsymbol{w}_{t},\mathcal{D}i)$
7:     end for
8:     $\boldsymbol{w}_{t+1}\leftarrow\sum_{i\in\mathcal{S}}p^{\prime}_{i}\boldsymbol{w}_{i}^{t+1}$
9:  end for
10:  return $\boldsymbol{{w}}^{u}\leftarrow\boldsymbol{w}_{T}$
Federated Learning (FL). Suppose there is a client set $\mathcal{N}$ ($|\mathcal{N}|=N$), contributing to FL training. Each client $i\in\mathcal{N}$ has a local training dataset with size $n_{i}$, and the data is non-IID across different clients. The optimal FL model is defined as $\boldsymbol{w}^{*}=\arg\min_{\boldsymbol{w}}[F(\boldsymbol{w}):=\sum_{i\in\mathcal{N}}p_{i}f_{i}(\boldsymbol{w})]$, where $f_{i}$ denotes the local objective function of client $i$ with the aggregation weight $p_{i}=\frac{n_{i}}{\sum_{k\in\mathcal{N}}n_{k}}$. The empirically trained model $\boldsymbol{\boldsymbol{w}}^{o}$ is obtained by training all clients $i\in\mathcal{N}$, serving as the empirical approximation of the theoretical optimum $\boldsymbol{w}^{*}$ (summarized in Figure 3).
Federated Unlearning (FU). FU mechanisms aim to obtain an unlearned model $\boldsymbol{{w}}^{u}$ that removes the influence of client set $\mathcal{J\in\mathcal{N}}$ who requests to be forgotten from the original trained model $\boldsymbol{\boldsymbol{w}}^{o}$. The optimal unlearned model could be defined as $\boldsymbol{w}^{r*}=\arg\min_{\boldsymbol{w}}[F_{-\mathcal{J}}(\boldsymbol{w}):=\sum_{i\notin\mathcal{J}}p_{i}^{\prime}f_{i}(\boldsymbol{w})]$, where $p_{i}^{\prime}$ is the normalized aggregation weight exclude unlearned clients, i.e., $p_{i}^{\prime}=\frac{p_{i}}{1-P_{\mathcal{J}}}$ with $P_{\mathcal{J}}=\sum_{j\in\mathcal{J}}p_{j}$. The exact retrained model $\boldsymbol{w}^{r}$ is obtained after retraining all remaining clients [^2] In this paper, the retraining employs FedAvg [35]., serving as the empirical approximation of the theoretical optimum $\boldsymbol{w}^{r*}$ (summarized in Figure 3).
Setting. This paper focuses on an FU mechanism that starts with the original trained model $\boldsymbol{\boldsymbol{w}}^{o}$ and iteratively updates on the remaining clients’ data ($i\notin\mathcal{J}$) as shown in Algorithm 1 [52]. The goal is to obtain an unlearned model $\boldsymbol{{w}}^{u}$ that closely replicates the results of exact retraining on the remaining data distribution. This approach aligns with the perspective of continual learning for unlearning [43, 30, 38].
@@ 4.1 FU Verification
FU verification is to evaluate how well the unlearning mechanism effectively removes the data [53]. Previous studies often employ the exact retrained model $\boldsymbol{w}^{r}$ as the evaluation benchmark [14, 6], but this benchmark suffers from variability due to randomness in the retraining process. Thus, we employ the optimal unlearned model $\boldsymbol{w}^{r*}$ as a theoretical benchmark, allowing for reliable FU evaluations in theoretical analysis. We introduce a verification metric as the performance disparity between the unlearned model and $\boldsymbol{w}^{r*}$:
@@
where $F_{-\mathcal{J}}(\cdot)$ measures performance over the remaining clients after unlearning clients set $\mathcal{J}$. The expectation $\mathbb{E}$ is taken over the randomness in obtaining $\boldsymbol{{w}}^{u}$ via $\mathcal{M}$.
@@ 4.2 Global Stability
Effective unlearning often requires modification to the trained model, which potentially alters global performance and compromises the utility of other remaining clients, especially under data heterogeneity. To measure the extent of performance stability in FU, we introduce the global stability metric as follows:
@@ 4.3 Local Fairness
In FL, fairness can be associated with the consistency of model performance across different clients. Specifically, a model is considered fairer if its performance has a smaller variance across clients [28].
@@
where $\Delta f_{i}(\boldsymbol{{w}}^{u})=\mathbb{E}\left[f_{i}(\boldsymbol{{w}}^{u})\right]-f_{i}(\boldsymbol{w}^{*})$ represents the utility change for remaining client $i\notin\mathcal{J}$.
The metric $Q$ captures FU’s impact on utility changes experienced by remaining clients and further facilitates theoretical analysis of fairness implication in Section 5.2.
@@ 5 Theoretical Analysis on Trade-offs in FU
This section provides a theoretical analysis of the trade-offs in FU, particularly focusing on the balance between FU verification and stability, as well as FU verification and fairness, as outlined in Section 5.1 and 5.2, respectively. Our analyses underscore the inevitable costs and trade-offs that any practical FU mechanism must address, particularly in the presence of data heterogeneity. By identifying these trade-offs, we provide a principled foundation for developing heterogeneity-aware FU algorithms that manage the trade-offs. To begin, we formally state the assumptions required for the theoretical analysis.
@@
5.1 assumes the data heterogeneity with parameters $\zeta_{\mathcal{S}},\beta_{\mathcal{S}}$, representing the degrees of data heterogeneity within the selected subset $\mathcal{S}$ of remaining clients, aligning closely with the framework presented by Wang et al. [47].
@@
Assume that local objective functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are all $\mu$-strong convex. For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, $f_{i}$ satisfies the following inequality: $f_{i}(\mathbf{u})\geq f_{i}(\mathbf{v})+\langle\nabla f_{i}(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle+\frac{\mu}{2}\|\mathbf{u}-\mathbf{v}\|^{2}$, where $\mu>0$ is the convexity constant.
@@
Assume that local objective functions $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are all $L$-smooth. For any vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{d}$, $f_{i}$ satisfies the following inequality: $f_{i}(\mathbf{u})\leq f_{i}(\mathbf{v})+\langle\nabla f_{i}(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle+\frac{L}{2}\|\mathbf{u}-\mathbf{v}\|^{2}$, where $L>0$ is the Lipschitz constant of the gradient of $f_{i}$.
@@
Let $\xi_{t}^{k}$ be sampled from the $k$-th client’s local data uniformly at random. The variance of stochastic gradients in each client is bounded at round $t$: $\mathbb{E}\left\|\boldsymbol{g}_{k}(\boldsymbol{w}_{t})-\nabla f_{k}\left(\boldsymbol{w}_{t}\right)\right\|^{2}\leq\sigma_{k,t}^{2}$ for $k=1,\cdots,N$, where $\boldsymbol{g}_{k}(\boldsymbol{w}_{t})=\nabla f_{k}\left(\mathbf{w}_{t},\xi_{t}\right)$.
@@
5.2-5.4 are commonly used for the FL convergence analysis, e.g., Li et al. [29], Wang et al. [47] and 5.5 assumes unlearned clients’ aggregate weights do not exceed those of the remaining clients. This is crucial to prevent catastrophic consequences, which could undermine the objectives of fairness and stability in FU.